# EPiC: Cost-effective Search-based Prompt Engineering of LLMs for Code Generation

### v1

@@ 3.1 Initial Evaluation
3.2 Evolutionary Prompt Engineering
In the EPE phase, the first step (Step Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
3
) involves populating the first generation of individuals, where each individual is a distinct prompt to generate the same code. In Step Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
3
@@ 3.2 Evolutionary Prompt Engineering
, we generate multiple prompts by modifying the initial prompt using an LLM agent. Generating this prompt population forms an important part of our evolutionary algorithm. A more detailed description of this process will be provided in Section 3.2.1. This step corresponds to line 8 of Algorithm 1.
In Step Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
4
@@
of EPE which corresponds to lines 12 to 20 in Algorithm 1, we generate code for each prompt and evaluate each generated code sample using the same test cases. We define a fitness function based on the ratio of test cases passed. If any of the prompts achieves the maximum fitness score, the process stops. If not, the process continues until all prompts are validated and given a fitness score, which will guide the selection process. The fitness function is basically the passing rate of each prompt. The mathematical formula is explained in Section 4.4.
This fitness function is used to select the candidate prompts for mutation in Step Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
5
@@
. This is implemented in chooseCandidate function in Algorithm 1 (Lines 22 and 26), where a weighted random selection algorithm selects candidates randomly, with the probability of selection being proportional to their respective fitness score (Section 4.4). This selection occurs with substitution, allowing the same prompt to be chosen multiple times. In Step Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
5
, we select $N-1$ candidates for mutation and one prompt for “elitism”.
Following prompt selection, in Step Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
6
@@
, EPiC randomly mutates the selected prompt candidates to better explore the search space of potential prompts in the next generation. The prompt mutation which is detailed in Algorithm 2, is elaborated in Section 3.2.2. The next generation is formed by adding the elite prompt to the pool of mutated prompts. Lines 21 to 27 in Algorithm 1 corresponds to Steps Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
5
and Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
6
@@
Figure 4: An example of a mutation generated by $sim\_words\_as\_mutator$. The orange words highlight the selected and mutated words in this example.
In the next two subsections, we will explain more details about Step Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
3
(Initial Population Builder), and Step Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
6
(Prompt Mutation.)
3.2.1 Initial Population Builder
This subsection provides a detailed explanation of the Step Test Case Generation Prompt
You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:
def add3Numbers(x, y, z):
unit tests:
assert add3Numbers(1, 2, 3) == 6
assert add3Numbers(-1, 2, 3) == 4
assert add3Numbers(1, -2, 3) == 2
assert add3Numbers(1, 2, -3) == 0
assert add3Numbers(-3, -2, -1) == -6
assert add3Numbers(0, 0, 0) == 0
3
in Figure 2. Human-written prompts are often very brief and lack the necessary information or context. This results in insufficient descriptions and a failure to adhere to a well-defined format [43]. Thus manual prompting typically results in several back-and-forth interactions with LLM until all necessary details are given [28]. Providing the initial prompt in a structured and elaborate format will guide the LLM to generate better results with fewer interactions [31].
In our evolutionary algorithm (Algorithm 1), we require an initial population of prompts. To create it, we utilize OpenAI’s GPT-4o to generate multiple prompts based on the initial given prompt. The prompt we use is as follows:
Please rewrite the function description based on these instructions:
1- Add input and output types of the function to the description.
2- Elaborate the description so that it is understandable for large language models. 3- Keep the original test cases and add three test cases to the description to cover the edge cases. Do not separate the generated test cases and the original ones. 4- Keep the structure of the function and add the description as a comment in the function. Use at most 500 words.
@@ 3.2.1 Initial Population Builder
Figure 3(a) illustrates the original prompt provided to an LLM, i.e., GPT-4o, which results in an implementation that fails to pass the test cases. Conversely, Figure 3(b) depicts a transformed version of the prompt that successfully guides the LLM towards a correct implementation. The primary issue with the original implementation is the incorrect initialization of the Eulerian number. In contrast, the transformed prompt includes specific instructions on how to properly initialize the Eulerian number, leading to a correct implementation. The orange text in both figures highlights the code and text related to the initialization.
After generating the elaborate prompt, we add the previously generated test cases (Step Please rewrite the function description based on these instructions:
1- Add input and output types of the function to the description. 2- Elaborate the description so that it is understandable for large language models. 3- Keep the original test cases and add three test cases to the description to cover the edge cases. Do not separate the generated test cases and the original ones. 4- Keep the structure of the function and add the description as a comment in the function. Use at most 500 words.
1
@@
in Figure 2) to each prompt. These test cases will also be used during the evolutionary algorithm’s fitness evaluation in Step Please rewrite the function description based on these instructions:
1- Add input and output types of the function to the description. 2- Elaborate the description so that it is understandable for large language models. 3- Keep the original test cases and add three test cases to the description to cover the edge cases. Do not separate the generated test cases and the original ones. 4- Keep the structure of the function and add the description as a comment in the function. Use at most 500 words.
4
3.2.2 Prompt Mutation
This subsection provides a detailed explanation of the Step Please rewrite the function description based on these instructions:
1- Add input and output types of the function to the description. 2- Elaborate the description so that it is understandable for large language models. 3- Keep the original test cases and add three test cases to the description to cover the edge cases. Do not separate the generated test cases and the original ones. 4- Keep the structure of the function and add the description as a comment in the function. Use at most 500 words.
6
@@ 3.2.2 Prompt Mutation
in Figure 2. We employed two kinds of prompt mutation approaches. In the first approach, EPiC uses an LLM to facilitate the mutation process ($LLM\_as\_mutator$). This approach is inspired by a framework named EvoPrompt [16]. In $LLM\_as\_mutator$, we provide LLMs with predefined instructions on how to implement the prompt mutation. In the second approach, we employ Natural Language Processing (NLP) techniques using the NLTK and Gensim libraries to alter the prompts by substituting words with their synonyms ($sim\_words\_as\_mutator$). This method is our contribution as a local lightweight prompt mutation process. The main motivator is to reduce the cost of the iterative part of the EA. This prompt mutation algorithm ($sim\_words\_as\_mutator$) is described in Algorithm 2. We will compare the performance of both approaches in the following sections. We use the following prompt for mutation using LLMs:
@@
You are a mutation tool. This is a Python function and its description. Please mutate the description by enhancing its clarity and comprehension for sophisticated language models.
Please put the changed description between #Explanation and #End. Use at most 600 words.
@@
In the $sim\_words\_as\_mutator$ algorithm (Algorithm 2), we first extract the description part of the prompt. For each word in the sentence, if the word is not a stop word or proper noun, it is lemmatized, and the get_related_words() function (line 10 in Algorithm 2) is called to get its related words. To obtain the related words for a given $word$, we use $WordNet$ and $gensim$’s $GloVevectors$ ($fasttext-wiki-news-subwords-300$) to calculate the cosine similarity between the word and other related words. We rank the related words to create a list of the most similar words to the original word, and based on the specified $num\_t$ or $sim\_t$, we filter out the least related words and keep words that pass the given threshold. Then, we randomly substitute the word with one of the related words based on a $mutation\_probability$. Figure 4 illustrates an example of mutation performed by $sim\_words\_as\_mutator$. In this example, selected words are highlighted and replaced with their related synonyms by this algorithm. For instance, the word “takes” is randomly chosen for mutation. The algorithm first identifies similar words for this word, resulting in the following list: [“take”, “make”, “require”, “have”, “carry”, “get”, “bring”, “accept”, “lead”, “hold”]. The algorithm then randomly selects the word “accept” and replaces it with “take”. This process is similarly applied to other randomly selected words, such as “iterates”, “returns”, “containing”, “collection”, and “elements”.
4 Experiment Design
4.1 Research Questions
We define the following research questions to explore the performance of EPiC:
  • RQ1: How cost-effective are the SOTA LLM-based code generation approaches? Motivation: Most of the SOTA LLM-based code generation studies do not include a comprehensive approach to evaluate cost-effectiveness. In this RQ, we set out to evaluate these tools from the cost-effectiveness perspective.
  • RQ2: How effective is EPiC in code generation? Motivation: In this RQ, we explore the performance of EPiC against SOTA LLM-based code generation tools in terms of both cost and effectiveness. We experiment with a large closed-source LLM (RQ2.1) and a smaller open-source model(RQ2.2)
    • RQ2.1: How cost-effective is EPiC compared to the SOTA?
    • RQ2.2: How effective is EPiC on smaller open-source LLMs compared to larger closed-source models?
  • RQ3:What are the impacts of different EA design choices on the cost-effectiveness of EPiC? Motivation: The goal of this RQ is to explore the impact of different hyperparameters in our approach. We define three sub-RQs to address the impact of the mutation generator (RQ3.1), mutation probability and similar word selection (RQ3.2), and population size (RQ3.3), as follows:
    • RQ3.1: How does the complexity of the mutation generator impact the cost-effectiveness of EPiC?
    • RQ3.2: How does the mutation probability and similar word selection impact the performance of EPiC?
    • RQ3.3: How does population size impact the cost-effectiveness of EPiC?
4.2 Datasets and Models
TABLE I: Default configuration parameters
$mutation\_probability$ $population\_size$ $top\_n$ $mutation\_alg$ $llm$
$0.4$ $5$ $10$ $sim\_words\_as\_mutator$ $gpt4o$
@@ 4.2 Datasets and Models
We used two datasets for our experiments: HumanEval [8] and MBPP [24].
The HumanEval is a collection of 164 programming problems designed to evaluate the functional correctness of code generated by AI models. Each problem includes a function signature, a detailed description of the task (docstring), the function body, and multiple unit tests to verify the solution. One major obstacle in making fair judgments on the performance of the related works was that some papers did not use the HumanEval dataset as is. For instance, there are cases where the samples are removed or changed. To have a fair comparison among all methods we transformed the original dataset into each tool’s required format (if needed) to re-run their code using the original dataset. We will explain this in more detail as we discuss the configuration setup for assessing each related work.
The Mostly Basic Programming Problems (MBPP) comprises 974 short Python programming tasks designed to be solvable by entry-level programmers. The tasks were crowd-sourced from individuals with basic Python knowledge, who provided a problem statement, a self-contained Python function solving the problem, and three test cases to check for correctness. A subset of 427 tasks was manually inspected and edited for consistency and accuracy, referred to as mbpp-sanitized. In this paper, we used mbpp-sanitized for our experiments. While experimenting with the SOTA, we discovered that they have used different subsets of the MBPP dataset. For this reason, their reported results are not comparable. Consequently, we chose to use the mbpp-sanitized subset in all our experiments.
For RQ1 and RQ2.1, we used both datasets. For RQ2.2 and RQ3, we used only HumanEval. We will explain this in more detail in the corresponding RQs.
We employed two LLMs in our study: OpenAI’s GPT-4o [^2] https://openai.com/index/hello-gpt-4o/ and Magicoder-S-DS-6.7B [11]. The latter model was specifically utilized for RQ 2.2, while OpenAI’s GPT-4o model was used for all other experiments.
4.3 Evaluation Metrics

### v2

@@ 3.1. Initial Evaluation
Code Generation Prompt You are a Python developer who implements the correct code based on the function description provided. You are given one or more functions to implement. Do not delete import statements from the code snippet. Use at most 1000 words.
3.2. Evolutionary Prompt Engineering
@@
Evolutionary Prompt Engineering (EPE) is a systematic approach that leverages EA to iteratively refine and optimize prompts for code generation. In the EPE phase, the first step (Step 3) involves populating the first generation of individuals, where each individual is a distinct prompt to generate the same code. In Step 3, we generate multiple prompts by modifying the initial prompt using an LLM agent. Generating this prompt population forms an important part of our evolutionary algorithm. A more detailed description of this process will be provided in Section 3.2.1. This step corresponds to line 8 of Algorithm 1.
@@
In Step 4 of EPE, which corresponds to lines 12 to 20 in Algorithm 1, we generate code for each prompt and evaluate each generated code sample using the same test cases. We define a fitness function based on the ratio of test cases passed. If any of the prompts achieves the maximum fitness score, the process stops. If not, the process continues until all prompts are validated and given a fitness score, which will guide the selection process. The fitness function is the passing rate of each prompt. The mathematical formula is explained in Section 4.4.
@@
This fitness function is used to select the candidate prompts for mutation in Step 5. This is implemented in chooseCandidate function in Algorithm 1 (Lines 22 and 26), where a weighted random selection algorithm selects candidates randomly, with the probability of selection being proportional to their respective fitness score (Section 4.4). This selection occurs with substitution, allowing the same prompt to be chosen multiple times. In Step 5, we select $N-1$ candidates for mutation and one prompt for “elitism”.
@@
Following prompt selection, in Step 6, EPiC randomly mutates the selected prompt candidates to better explore the search space of potential prompts in the next generation. The prompt mutation which is detailed in Algorithm 2, is elaborated in Section 3.2.2. The next generation is formed by adding the elite prompt to the pool of mutated prompts. Lines 21 to 27 in Algorithm 1 correspond to Steps 5 and 6.
@@
In the next two subsections, we will explain the details of Step 3 (Initial Population Builder) and Step 6 (Prompt Mutation).
(a) Original prompt and its result
(b) Transformed prompt and its result
Figure 3. An unsuccessful original prompt on the left and the transformed successful version of it on the right. The main issue with the original implementation is the incorrect initialization of the Eulerian number. The orange text in both figures highlights the code and text related to the initialization of the function.
3.2.1. Initial Population Builder
This subsection provides a detailed explanation of Step 3 in Figure 2. Human-written prompts are often very brief and lack the necessary information or context. This results in insufficient descriptions and a failure to adhere to a well-defined format (Mu et al., 2023). Thus, manual prompting typically results in several back-and-forth interactions with LLM until all necessary details are given (Shin et al., 2023). Providing the initial prompt in a structured and elaborate format will guide the LLM to generate better results with fewer interactions (Jin et al., 2024).
In our evolutionary algorithm (Algorithm 1), we require an initial population of prompts. We utilize the LLM to augment multiple prompts based on the initial given prompt. The prompt we use is as follows:
Please rewrite the function description following the instructions: 1- Add input and output types of the function to the description. 2- Elaborate the description so that it is understandable for large language models. 3- Keep the original test cases and add three test cases to the description to cover the edge cases. Do not separate the generated test cases and the original ones. 4- Keep the structure of the function and add the description as a comment in the function. Use at most 500 words.
@@ 3.2.1. Initial Population Builder
Figure 3(a) illustrates the original prompt provided to an LLM, i.e., GPT-4o, which results in an implementation that fails to pass the test cases. Conversely, Figure 3(b) depicts a transformed version of the prompt that successfully guides the LLM toward the correct implementation. The primary issue with the original implementation is the incorrect initialization of the Eulerian number. In contrast, the transformed prompt includes specific instructions on how to properly initialize the Eulerian number, leading to a correct implementation. The orange text in both figures highlights the code and text related to the initialization.
@@
After generating the elaborate prompt, we add the previously generated test cases (Step 1 in Figure 2) to each prompt. These test cases will also be used during the evolutionary algorithm’s fitness evaluation in Step 4.
Figure 4. An example of a mutation generated by $sim\_words\_as\_mutator$. The orange words highlight the selected and mutated words in this example.
3.2.2. Prompt Mutation
@@
This subsection provides a detailed explanation of Step 6 in Figure 2. We employed two kinds of prompt mutation approaches. In the first approach, EPiC uses an LLM to facilitate the mutation process ($LLM\_as\_mutator$). This approach is inspired by a framework named EvoPrompt (Yang et al., 2023). In $LLM\_as\_mutator$, we provide LLMs with predefined instructions on how to implement the prompt mutation. In the second approach, we employ Natural Language Processing (NLP) techniques using the NLTK and Gensim libraries to alter the prompts by substituting words with their synonyms ($sim\_words\_as\_mutator$). Synonym replacement, when applied with controlled probability, can improve model performance without altering the fundamental semantics of a sentence (Wei and Zou, 2019). We extend this concept by integrating $sim\_words\_as\_mutator$ into our prompt evolution framework, reducing computational cost while making meaningful mutations. Since the evolutionary algorithm involves iterative prompt refinement, the cost-effectiveness and linguistic validity of this approach make it an ideal mutation operator. This prompt mutation algorithm ($sim\_words\_as\_mutator$) is described in Algorithm 2. We will compare the performance of both approaches in Section 6. We use the following prompt for the mutation process in $LLM\_as\_mutator$:
@@
You are a mutation tool. This is a Python function and its description. Please change the description by enhancing its clarity and comprehensibility for sophisticated language models. Please put the changed description between #Explanation and #End. Use at most 600 words.
@@
In the $sim\_words\_as\_mutator$ (Algorithm 2), we first extract the description part of the prompt. For each word in the sentence, if the word is not a stop word or proper noun, it is lemmatized, and the get_related_words() function (line 10 in Algorithm 2) is called to get its related words. To obtain the related words for a given $word$, we use $WordNet$ and $gensim$’s $GloVevectors$ ($fasttext-wiki-news-subwords-300$) to calculate the cosine similarity between the word and other related words. We rank the related words to create a list of the most similar words to the original word, and based on the specified $num\_t$ or $sim\_t$, we filter out the least related words and keep words that pass the given threshold. Then, we randomly substitute the word with one of the related words based on a $mutation\_probability$. Figure 4 illustrates an example of mutation performed by $sim\_words\_as\_mutator$. In this example, selected words are highlighted and replaced with their related synonyms by this algorithm. For instance, the word “takes” is randomly chosen for mutation. The algorithm first identifies similar words for this word, resulting in the following list: [“take”, “make”, “require”, “have”, “carry”, “get”, “bring”, “accept”, “lead”, “hold”]. The algorithm then randomly selects the word “accept” and replaces it with “take”. This process is similarly applied to other randomly selected words, such as “iterates”, “returns”, “containing”, “collection”, and “elements”.
4. Experiment Design
4.1. Research Questions
We define the following research questions to evaluate the performance of EPiC:
RQ1: How does EPiC perform across different SOTA LLMs and datasets in terms of pass@1 and token usage? Motivation: Evaluating EPiC on the SOTA LLMs provides us with insights into their performance and token usage in code generation. This evaluation also helps us find the best-performing LLM for the next RQ. In addition, it also shows that EPiC’s performance is not limited to one LLM.
RQ2: How does EPiC compare to other iterative-based agents (Reflexion, LATS, and LDB) in terms of pass@1, token usage, and cost-effectiveness across different benchmarks? Motivation: cost and pass@1 are the two major factors in identifying the most effective agent. In this RQ, we analyze the cost-effectiveness of the SOTA code generation agents and compare them with EPiC.
4.2. Datasets and Models
@@
We used three datasets for our experiments: HumanEval+ (Liu et al., 2023a), MBPP+ (Liu et al., 2023a), and BigCodeBenchHard (Zhuo et al., 2024).
HumanEval+ (Chen et al., 2021) is a collection of 164 programming problems designed to evaluate the functional correctness of code generated by AI models. Each problem includes a function signature, a detailed description of the task (docstring), the function body, and multiple unit tests to verify the solution. HumanEval+ extends the original HumanEval dataset by increasing the number of test cases by 80 times using both LLM-based and mutation-based automated test generation techniques. This augmentation helps uncover incorrect LLM-generated code that was previously misclassified as correct due to inadequate testing, leading to more accurate pass@k performance evaluations for LLMs.
MBPP+ (Liu et al., 2023a) builds upon the MBPP (Mostly Basic Python Programming) dataset, which consists of 974 crowd-sourced programming problems, each with a problem statement, function signature, and three test cases. MBPP+ enhances this dataset by incorporating additional, automatically generated test cases to improve the robustness of functional correctness evaluations for LLMs. Similar to HumanEval+, this augmentation ensures that LLM-generated code is tested against a wider variety of scenarios, including edge cases, leading to more reliable performance assessments.
BigCodeBenchHard (Zhuo et al., 2024) is a challenging subset of BigCodeBench designed to evaluate LLMs on complex programming tasks requiring diverse function calls, compositional reasoning, and precise instruction following. Covering 139 libraries across seven domains, it includes rigorous execution-based evaluation with an average of 5.6 test cases per task. Compared to benchmarks such as HumanEval, it presents longer prompts, higher cyclomatic complexity, and broader tool use.
We employed three SOTA LLMs from different vendors in our study: OpenAI’s O3-mini [^2] https://openai.com/index/openai-o3-mini/, Antropic’s Claude 3.7 Sonnet [^3] https://www.anthropic.com/claude/sonnet, and Deepseek V3444https://github.com/deepseek-ai/DeepSeek-V3.
4.3. Evaluation Metrics
4.3.1. pass@k